# On the structure of length sets with maximal elasticity

### v1

@@ 1. Introduction
Let $H$ be a Krull monoid or a Krull domain with class group $G$ and suppose that each class contains a prime divisor. Then each (non-zero and non-invertible) element $a\in H$ has a factorization into atoms, say $a=u_{1}\cdot\ldots\cdot u_{k}$, where $k$ is the factorization length and $u_{1},\ldots,u_{k}$ are atoms (irreducible elements) of $H$. The set $\mathsf{L}(a)\subset\mathbb{N}$ of all possible factorization lengths $k$ is called the length set of $a$. Length sets and all parameters controlling their structure (such as elasticities and sets of distances) are the best investigated invariants in factorization theory (for a survey, we refer to [20]). To begin with the extremal cases, we have $|\mathsf{L}(a)|=1$ for all relevant $a\in H$ if and only if $|G|\leq 2$. If $G$ is infinite, then every finite nonempty subset of $\mathbb{N}_{\geq 2}$ occurs as a length set of some $a\in H$. Now suppose that $G$ is finite with $|G|\geq 3$. Then the length sets can be arbitrarily large, but they are all finite and well-structured (Proposition 2.2). Roughly speaking, they are sort of generalized arithmetic progressions. In addition, it is known that almost all length sets are intervals (Proposition 3.1).
For a finite nonempty set $L\subset\mathbb{N}$, we denote by $\rho(L)=\max L/\min L$ the elasticity of $L$, and by $\rho(H)$, defined as the supremum of $\rho\big{(}\mathsf{L}(a)\big{)}$ over all $a$ of $H$, the elasticity of $H$. It is well-known that $\rho(H)=\mathsf{D}(G)/2$, where $\mathsf{D}(G)$ denotes the Davenport constant of $G$ (see Subsection 2.2). Recent research has seen much interest on the structure of length sets with maximal elasticity. It has turned out that these length sets have a much simpler structure than length sets in general (compare Propositions 2.2 and 2.4).
@@ 2. Preliminaries
We denote by $\mathbb{N}$ the set of positive integers and by $\mathbb{N}_{0}$ the set of non-negative integers. For real numbers $a,b\in\mathbb{R}$, $a,b]=\{x\in\mathbb{Z}:a\leq x\leq b\}$ means the discrete interval between $a$ and $b$. Let $L,L^{\prime}\subset\mathbb{Z}$ be subsets. Then $L+L^{\prime}=\{a+b:a\in L,\hskip 2.84544ptb\in L^{\prime}\}$ denotes their sumset. The set of distances $\Delta(L)\subset\mathbb{N}$ is the set of all $d\in\mathbb{N}$ for which there is $a\in L$ such that $L\cap[a,a+d]=\{a,a+d\}$. If $L^{\prime}\subset L$ is finite, then $L^{\prime}$ is said to be an interval of $L$ if $L^{\prime}\neq\emptyset$ and $L^{\prime}=L\cap[\min L^{\prime},\max L^{\prime}]$. By an interval, we mean an interval of $\mathbb{Z}$, whence an interval is an arithmetic progression with difference $1$.
Let $G$ be an additively written finite abelian group. Let $r\in\mathbb{N}$ and $(e_{1},e_{2},\ldots,e_{r})$ be an $r$-tuple of non-zero elements of $G$, then $(e_{1},e_{2},\ldots,e_{r})$ is said to be independent if for all $(m_{1},m_{2},\ldots,m_{r})\in\mathbb{Z}^{r}$ an equation $m_{1}e_{1}+m_{2}e_{2}+\ldots+e_{r}m_{r}=0$ implies $m_{i}e_{i}=0$ for all $i\in[1,r]$. An independent $r$-tuple of non-zero elements $(e_{1},e_{2},\ldots,e_{r})$ is said to be a basis if $G=\langle e_{1}\rangle\oplus\langle e_{2}\rangle\oplus\ldots\oplus\langle e_{r}\rangle$. For $n\in\mathbb{N}$, we let $C_{n}$ denote a cyclic group of order $n$.
@@
Let $H$ be an atomic monoid and suppose that $x,y\in\mathsf{Z}(H)$ and $z=\textup{gcd}(x,y)$. Then we call $\mathsf{d}(x,y)=\textup{max}\{|z^{-1}x|,|z^{-1}y|\}$ the distance between $x$ and $y$. The distance function $\mathsf{d}:\mathsf{Z}(H)\times\mathsf{Z}(H)\longrightarrow\mathbb{N}_{0}$ is a metric. We define the catenary degree $\mathsf{c}(a)$ for $a\in H$ to be the smallest $N\in\mathbb{N}_{0}\cup\{\infty\}$ such that, for any two factorizations $z,z^{\prime}\in\mathsf{Z}(a)$, there exists a finite sequence $z=z_{0},z_{1},\ldots,z_{k}=z^{\prime}\in\mathsf{Z}(a)$ satisfying that $\mathsf{d}(z_{i-1},z_{i})\leq N$ for all $i\in[1,k]$. If $z,\,z^{\prime}\in\mathsf{Z}(a)$ and $z\neq z^{\prime}$, then
@@ 2.2. Krull Monoids
    There is a divisor homomorphism $\varphi\colon H\to D$, where $D$ is a free abelian monoid such that for every $\alpha\in D$ there is a finite nonempty set $A\subset H$ such that $\alpha=\gcd\varphi(A)$.
@@
If $D$ is a domain, then $D$ is a Krull domain if and only if its monoid of nonzero elements is a Krull monoid. Condition (a) shows that every integrally closed noetherian domain is Krull. For background and further examples of Krull monoids, we refer to [11, 6].
@@
Furthermore, $\mathcal{B}(G_{0})=\{S\in\mathcal{F}(G_{0}):\sigma(S)=0\}$ is the monoid of zero-sum sequences over $G_{0}$. $\mathcal{B}(G_{0})$ is a finitely generated monoid, and since the inclusion $\mathcal{B}(G_{0})\hookrightarrow\mathcal{F}(G_{0})$ is a divisor homomorphism, Condition (c) implies that $\mathcal{B}(G_{0})$ is a Krull monoid. We denote $\mathcal{A}(G_{0})$ the set of atoms of $\mathcal{B}(G_{0})$, and we call the elements of $\mathcal{A}(G_{0})$ the minimal zero-sum sequences over $G_{0}$. We set
@@
Lastly, the Davenport constant of a finite abelian group $G$ is defined by
@@
A transfer homomorphism $\theta:H\longrightarrow B$ allows us to pull back arithmetic properties from $B$ to $H$. In particular, we have
@@ 2.4. Length Sets with Maximal Elasticity
Our goal is to understand the structure of length sets with maximal elasticity. To begin with, we recall the Structure Theorem for Length Sets (this result will be needed later on in the proof of Theorem 1.2). To do so, we need the concepts of AAMPs (almost arithmetic multiprogressions) and of AAPs (almost arithmetic progressions).
@@
  • $\min L^{*}=0$, $L^{*}$ is an interval of $\min L^{*}+\mathcal{D}+d\mathbb{Z}$, $\ell$ is maximal with $\min L^{*}+(\ell-1)d\in L^{*}$,
@@
A realization theorem by Schmid shows that the above structural description is best possible ([18]). In order to describe the structure of length sets with maximal elasticity, we need one more invariant and then we state some results as well as a standing conjecture (Conjecture 2.3) about length sets with maximal elasticity. Let $\Delta_{\rho}(G)$ denote the set of all $d\in\mathbb{N}$ with the following property:
@@
If $|G|>4$, then $\Delta_{\rho}(G)=\{1\}$ if and only if $G$ is neither cyclic nor an elementary $2$-group.
@@
Proposition 2.4.
Suppose $\Delta_{\rho}(G)=\{1\}$. There exists some $M\in\mathbb{N}_{0}$ such that every $L\in\mathcal{L}(H)$ with $\rho(L)=\mathsf{D}(G)/2$ is an AAP with difference $1$ and bound $M$.
If the above statement holds for $M$, then it holds for every $M^{\prime}$ with $M^{\prime}\geq M$. However, apart from some exceptional cases, the statement does not hold for $M=0$. Comparing Proposition 2.4 with Proposition 2.2, we observe that the structure of length sets with maximal elasticity is much more rigid than those of general length sets. Moreover, if $G$ be cyclic or an elementary $2$-group with $|G|>4$, then $\Delta_{\rho}(G)\neq\{1\}$ and the statement of Proposition 2.4 does not hold.
@@
For the proof of Theorem 1.1, we proceed in a series of lemmas.
@@
Using the inclusion $\mathsf{L}(a)+\mathsf{L}(b)\subset\mathsf{L}(ab)$ for all $a,b\in H$, we get $\min\mathsf{L}(ab)\leq\min\mathsf{L}(a)+\min\mathsf{L}(b)\leq\max\mathsf{L}(a)+\max\mathsf{L}(b)\leq\max\mathsf{L}(ab)$ and hence
@@
which implies that $|U_{i}|=\mathsf{D}(G)$ for all $i\in[1,k]$ and $|V_{j}|=2$ for all $j\in[1,l]$. The reverse implication is straightforward (compare also with [8][Lemma 3.2 (i)]). ∎
@@
Since $\mathcal{B}(G)$ is finitely generated, $\mathcal{A}(G)$ is finite, say $\{U\in\mathcal{A}(G):|U|=\mathsf{D}(G)\}=\{U_{1},\ldots,U_{\lambda}\}$. Let
@@
Let $H$ be a transfer Krull monoid over a finite abelian group $G$ with $|G|\geq 3$ and which satisfies Property P. By (2.2), it is sufficient to prove the assertion for the monoid $H=\mathcal{B}(G)$. Let $g_{1},g_{2}\in G$ and let $U_{1},U_{2}\in\mathcal{A}(G)$ such that
@@
As applications of Theorem 1.1, we present two corollaries, which demonstrate that almost all (in a combinatorial sense) length sets with maximal elasticity are intervals. To establish the first result, we require a lemma concerning the growth behavior of the $n$-fold sumset of a finite set. Khvanskii first proved that $n$-fold sumset of a finite set $A\subset\mathbb{Z}^{d}$ grows like a polynomial. His result was generalized, refined, and the bounds were made effective by many authors ([15, 10, 5]). We cite a version due to Granville, Shakan, and Walker ([10]).
@@
We have $\mathcal{A}(\mathcal{B}_{\rho}(G))\subset\mathcal{B}_{\rho}(G)\subset\mathcal{F}(G)$ and set $\Omega:=\psi(\mathcal{A}(\mathcal{B}_{\rho}(G)))$. In particular, the monoid $\mathcal{B}_{\rho}(G)$ is isomorphic to the monoid $S:=[\Omega]$ generated by $\Omega$.
@@
Let $\alpha_{N}\in\mathbb{N}$ be such that $\alpha_{N}N\geq\alpha$, we claim that
@@
Hence the claim follows. Next, since $|\{B\in\mathcal{B}_{\rho}(G):|B|\leq n-\alpha\}|\geq|\{B\in\mathcal{B}_{\rho}(G):|B|\leq n-\alpha_{N}N\}|$, it suffices to show that
@@
Let $G=C_{n_{1}}\oplus\ldots\oplus C_{n_{r}}$ with $r\in\mathbb{N}_{0}$ and $1<n_{1}\,|\,\ldots\,|\,n_{r}$, and set $\mathsf{D}^{*}(G)=1+\sum_{i=1}^{r}(n_{i}-1)$. A simple example shows that
@@
is satisfied in each of the following cases.
@@
    $G$ is a cyclic group of odd order.
@@
    $G$ is not cyclic and $\mathsf{D}(G)=\mathsf{D}^{*}(G)$.
@@
    $G$ has odd order and there is some $U\in\mathcal{A}(G)$ of length $|U|=\mathsf{D}(G)$ that is not squarefree.
@@
(a) Suppose $G$ is cyclic of odd order $n\geq 3$. Let $g\in G$ be any generator. Clearly, $2g\in G$ is also a generator of the group, since the order of $G$ is odd. Now, consider $U_{1}=g^{n}$ and $U_{2}=(2g)^{n}$. Since both $U_{1}$ and $U_{2}$ are atoms with length $\mathsf{D}(G)$, our assumption holds.
(b) Suppose that $G\cong C_{n_{1}}\oplus\ldots\oplus C_{n_{r}}$ with $1<n_{1}\mid\ldots\mid n_{r}$ and that $\mathsf{D}^{*}(G)=\mathsf{D}(G)$. Let $(e_{1},\ldots,e_{r})$ be a basis of $G$ with $\operatorname{ord}(e_{i})=n_{i}$ for $i\in[1,r]$ and with $r\geq 2$. Then the sequence $U=(e_{1}+e_{2}+\ldots+e_{r})\prod_{i=1}^{r}e_{i}^{n_{i}-1}$ has the property that $|U|=\mathsf{D}(G)$. Moreover, the map $\varphi:G\longrightarrow G$, given by $e_{i}\mapsto e_{i}$ for all $i\in[1,r-1]$ and $e_{r}\mapsto e_{r}+e_{r-1}$, is an isomorphism. Since $U$ is a minimal zero-sum sequence, so $\varphi(U)$ is too, and therefore, $|\varphi(U)|=|U|=\mathsf{D}(G)$ and $e_{r-1}+e_{r}\in\textup{supp }(\varphi(U))$. If we choose $U_{1}=U$ and $U_{2}=\varphi(U)$, then the condition holds.
(c) Suppose $G$ has odd order. Then the map defined by the multiplication by two, $\varphi:G\longrightarrow G,\hskip 5.69046ptg\mapsto 2g$ for every $g\in G$, is an isomorphism. Since $U$ is not squarefree, there exists $g\in G$ such that $g^{2}\hskip 2.84544pt|\hskip 2.84544ptU$. Then $\varphi(U)$ is a minimal zero-sum sequence of length $|\varphi(U)|=|U|=\mathsf{D}(G)$ and $2g\in\textup{supp }(\varphi(U))$. Thus, the condition holds with $U_{1}=U$ and $U_{2}=\varphi(U)$.
@@
We end this section by considering a Property P of finite abelian groups and its arithmetic consequences. Theorem 4.3 establishes a result on the catenary degree for elements with maximal elasticity. It is analogous to the result given by Proposition 3.1, but its validity is quite restricted. Let $H$ be an atomic monoid and let $a\in H$. Then the basic inequality (2.1) shows that $\mathsf{c}(a)\leq 3$ implies that $\mathsf{L}(a)$ is an interval. But the converse is far from being true. Thus, the statement in Theorem 4.3 is stronger than the statement in Theorem 1.1.
@@
Let $H$ be a transfer Krull monoid over a finite abelian group $G$ satisfying Property P. Then there exists some $a^{*}\in H$ such that all elements $a\in H$ with $\rho(\mathsf{L}(a))=\rho(H)$ have catenary degree $\mathsf{c}(aa^{*})\leq 3$, whence their length sets $\mathsf{L}(a^{*}a)$ are intervals with elasticity $\rho(H)$.
@@
where the sequences $A_{g}$ are as in the Definition of Property P. Let $A\in\mathcal{B}(G)$ be with $\rho(\mathsf{L}(A))=\mathsf{D}(G)/2$. Then $\mathsf{\rho}(\mathsf{L}(A^{*}A))=\mathsf{D}(G)/2$ by Lemma 3.2. Since $\textup{supp }(A^{*})\cup\{0\}=G$, we get $\textup{supp }(A^{*}A)\cup\{0\}=G$. Thus, [6, Theorem 7.6.8] implies that $\mathsf{c}(A^{*}A)\leq 3$, and consequently, the length set $\mathsf{L}(A^{*}A)$ forms an interval. It has elasticity $\rho(H)$ by Lemma 3.2. ∎

### v2

@@ 1. Introduction
Let $H$ be a Krull monoid or a Krull domain with class group $G$ and suppose that each class contains a prime divisor. Then each (non-zero and non-invertible) element $a\in H$ has a factorization into atoms, say $a=u_{1}\cdot\ldots\cdot u_{k}$, where $k$ is the factorization length and $u_{1},\ldots,u_{k}$ are atoms (irreducible elements) of $H$. The set $\mathsf{L}(a)\subset\mathbb{N}$ of all possible factorization lengths $k$ is called the length set of $a$. Length sets and all parameters controlling their structure (such as elasticities and sets of distances) are the best investigated invariants in factorization theory (for a survey, we refer to [20]). To begin with the extreme cases, we have $|\mathsf{L}(a)|=1$ for all relevant $a\in H$ if and only if $|G|\leq 2$. If $G$ is infinite, then every finite nonempty subset of $\mathbb{N}_{\geq 2}$ occurs as a length set of some $a\in H$. Now suppose that $G$ is finite with $|G|\geq 3$. Then the length sets can be arbitrarily large, but they are all finite and well-structured (Proposition 2.2). Roughly speaking, they are sorts of generalized arithmetic progressions. In addition, it is known that almost all length sets are intervals (Proposition 3.1).
For a finite nonempty set $L\subset\mathbb{N}$, we denote by $\rho(L)=\max L/\min L$ the elasticity of $L$, and by $\rho(H)$, defined as the supremum of $\rho\big(\mathsf{L}(a)\big)$ over all $a$ of $H$, the elasticity of $H$. It is well-known that $\rho(H)=\mathsf{D}(G)/2$, where $\mathsf{D}(G)$ denotes the Davenport constant of $G$ (see Subsection 2.2). Recent research has seen much interest in the structure of length sets with maximal elasticity. It has turned out that these length sets have a much simpler structure than length sets in general (compare Propositions 2.2 and 2.4).
@@ 2. Preliminaries
We denote by $\mathbb{N},\mathbb{N}_{0},$ and $\mathbb{P}$ the sets of positive integers, non-negative integers, and prime numbers, respectively. For real numbers $a,b\in\mathbb{R}$, $a,b]=\{x\in\mathbb{Z}:a\leq x\leq b\}$ means the discrete interval between $a$ and $b$. Let $L,L^{\prime}\subset\mathbb{Z}$ be subsets. Then $L+L^{\prime}=\{a+b:a\in L,\hskip 2.84544ptb\in L^{\prime}\}$ denotes their sumset. The set of distances $\Delta(L)\subset\mathbb{N}$ is the set of all $d\in\mathbb{N}$ for which there is $a\in L$ such that $L\cap[a,a+d]=\{a,a+d\}$. If $L^{\prime}\subset L$ is finite, then $L^{\prime}$ is said to be an interval of $L$ if $L^{\prime}\neq\emptyset$ and $L^{\prime}=L\cap[\min L^{\prime},\max L^{\prime}]$. By an interval, we mean an interval of $\mathbb{Z}$, whence an interval is an arithmetic progression with difference $1$.
Let $G$ be an additively written finite abelian group. Let $r\in\mathbb{N}$ and $(e_{1},e_{2},\ldots,e_{r})$ be an $r$-tuple of non-zero elements of $G$. Then $(e_{1},e_{2},\ldots,e_{r})$ is said to be independent if for all $(m_{1},m_{2},\ldots,m_{r})\in\mathbb{Z}^{r}$ an equation $m_{1}e_{1}+m_{2}e_{2}+\ldots+e_{r}m_{r}=0$ implies $m_{i}e_{i}=0$ for all $i\in[1,r]$. An independent $r$-tuple of non-zero elements $(e_{1},e_{2},\ldots,e_{r})$ is said to be a basis if $G=\langle e_{1}\rangle\oplus\langle e_{2}\rangle\oplus\ldots\oplus\langle e_{r}\rangle$. For $n\in\mathbb{N}$, we let $C_{n}$ denote a cyclic group of order $n$.
@@
Let $H$ be an atomic monoid and suppose that $x,y\in\mathsf{Z}(H)$ and $z=\textup{gcd}(x,y)$. Then we call $\mathsf{d}(x,y)=\textup{max}\{|z^{-1}x|,|z^{-1}y|\}$ the distance between $x$ and $y$. Explicitly, if $x=u_{1}\cdot\ldots\cdot u_{n}w_{1}\cdot\ldots\cdot w_{r}$ and $y=v_{1}\cdot\ldots\cdot v_{m}w_{1}\cdot\ldots\cdot w_{r}$ with $n,m,r\in\mathbb{N}_{0},$
$u_{1},\ldots,u_{n},v_{1},\ldots,v_{m},w_{1},\ldots,w_{r}\in\mathcal{A}(H_{\textup{red}})$ and $\{u_{1},\ldots,u_{n}\}\cap\{v_{1},\ldots,v_{m}\}=\emptyset$, then $\mathsf{d}(x,y)=\textup{max}\{n,m\}$. The distance function $\mathsf{d}:\mathsf{Z}(H)\times\mathsf{Z}(H)\longrightarrow\mathbb{N}_{0}$ is a metric. We define the catenary degree $\mathsf{c}(a)$ for $a\in H$ to be the smallest $N\in\mathbb{N}_{0}\cup\{\infty\}$ such that, for any two factorizations $z,z^{\prime}\in\mathsf{Z}(a)$, there exists a finite sequence $z=z_{0},z_{1},\ldots,z_{k}=z^{\prime}\in\mathsf{Z}(a)$ satisfying that $\mathsf{d}(z_{i-1},z_{i})\leq N$ for all $i\in[1,k]$. If $z,\,z^{\prime}\in\mathsf{Z}(a)$ and $z\neq z^{\prime}$, then
@@ 2.2. Krull Monoids
    There is a divisor homomorphism $\varphi\colon H\to D$, where $D$ is a free abelian monoid, such that for every $\alpha\in D$ there exists a nonempty finite set $A\subset H$ with $\alpha=\gcd\varphi(A)$.
@@
If $D$ is a domain, then $D$ is a Krull domain if and only if its monoid of nonzero elements is a Krull monoid. Condition (a) shows that every integrally closed noetherian domain is Krull. Let $H$ be a Krull monoid. Then there exists a free abelian monoid $F=\mathcal{F}(P)$ such that $H_{\textup{red}}\subset F$ is a submonoid, and the inclusion map $H_{\textup{red}}\hookrightarrow F$ is a divisor homomorphism that satisfies Condition (b). In this case, $F$ and $P$ are uniquely determined by $H$ (up to isomorphism). Consequently, the group
$\mathcal{C}(H)=\mathsf{q}(F)/\mathsf{q}(H_{\textup{red}}),$
where $\mathsf{q}(F)$ and $\mathsf{q}(H_{\textup{red}})$ are the quotient groups of $F$ and $H_{\textup{red}}$, respectively, depends only on $H$ and is called the class group of $H$. For background and further examples of Krull monoids, we refer to [11, 6].
@@
Furthermore, $\mathcal{B}(G_{0})=\{S\in\mathcal{F}(G_{0}):\sigma(S)=0\}$ is the monoid of zero-sum sequences over $G_{0}$. The monoid $\mathcal{B}(G_{0})$ is viewed as a submonoid of $\mathcal{F}(G_{0})$. $\mathcal{B}(G_{0})$ is a finitely generated monoid, and since the inclusion $\mathcal{B}(G_{0})\hookrightarrow\mathcal{F}(G_{0})$ is a divisor homomorphism, Condition (c) implies that $\mathcal{B}(G_{0})$ is a Krull monoid. We denote $\mathcal{A}(G_{0})$ the set of atoms of $\mathcal{B}(G_{0})$, and we call the elements of $\mathcal{A}(G_{0})$ the minimal zero-sum sequences over $G_{0}$. We set
@@
Lastly, the Davenport constant of $G$ is defined by
@@
It is known that $\rho(\mathcal{B}(G))=\mathsf{D}(G)/2$ for all finite abelian groups $G$.
@@
A transfer homomorphism $\theta:H\longrightarrow B$ allows us to pull back arithmetic properties from $B$ to $H$ ([6, Proposition 3.2.3]). In particular, we have
@@ 2.4. Length Sets with Maximal Elasticity
Our goal is to understand the structure of length sets with maximal elasticity. We begin by recalling the Structure Theorem for Length Sets (this result will be needed later on in the proof of Theorem 1.2). To do so, we need the concepts of AAMPs (almost arithmetic multiprogressions) and of AAPs (almost arithmetic progressions).
@@
  • $\min L^{*}=0$, $L^{*}$ is an interval of $\min L^{*}+\mathcal{D}+d\mathbb{Z}$, $\ell$ is maximal for the property that $\min L^{*}+(\ell-1)d\in L^{*}$,
@@
A realization theorem by Schmid shows that the above structural description is best possible ([18]). In order to describe the structure of length sets with maximal elasticity, we need one more invariant. Then we state some results as well as a standing conjecture (Conjecture 2.3) about length sets with maximal elasticity. Let $\Delta_{\rho}(G)$ denote the set of all $d\in\mathbb{N}$ with the following property:
@@
Let $G$ be a finite abelian group with $|G|>4$. Then $\Delta_{\rho}(G)=\{1\}$ if and only if $G$ is neither cyclic nor an elementary $2$-group.
@@
Proposition 2.4 ([2], Corollary 2.3).
Let $H$ be a transfer Krull monoid over a finite abelian group $G$ and suppose that $\Delta_{\rho}(G)=\{1\}$. Then there exists some $M\in\mathbb{N}_{0}$ such that every $L\in\mathcal{L}(H)$ with $\rho(L)=\mathsf{D}(G)/2$ is an AAP with difference $1$ and bound $M$.
If the above statement holds for $M$, then it holds for every $M^{\prime}$ with $M^{\prime}\geq M$. However, apart from some exceptional cases, the statement does not hold for $M=0$. Comparing Proposition 2.4 with Proposition 2.2, we observe that the structure of length sets with maximal elasticity is much more rigid than those of general length sets. Moreover, if $G$ is cyclic or an elementary $2$-group with $|G|>4$, then $\Delta_{\rho}(G)\neq\{1\}$ and the statement of Proposition 2.4 does not hold.
@@
For the proof of Theorem 1.1, we proceed via a series of lemmas.
@@
Using the inclusion $\mathsf{L}(a)+\mathsf{L}(b)\subset\mathsf{L}(ab)$ for all $a,b\in H$, we get
$\min\mathsf{L}(ab)\leq\min\mathsf{L}(a)+\min\mathsf{L}(b)\leq\max\mathsf{L}(a)+\max\mathsf{L}(b)\leq\max\mathsf{L}(ab)$
and hence
@@
which implies that $|U_{i}|=\mathsf{D}(G)$ for all $i\in[1,k]$ and $|V_{j}|=2$ for all $j\in[1,l]$. The reverse implication is straightforward. ∎
@@
Since $\mathcal{B}(G)$ is finitely generated, $\mathcal{A}(G)$ is finite, say
$\{U\in\mathcal{A}(G):|U|=\mathsf{D}(G)\}=\{U_{1},\ldots,U_{\lambda}\}.$
Let
@@
Let $H$ be a transfer Krull monoid over a finite abelian group $G$ with $|G|\geq 3$ and which satisfies Property P. By (2.2), it is sufficient to prove the assertion for the monoid $H=\mathcal{B}(G)$. Let $g_{1},g_{2}\in G$ and let $U_{1},U_{2}\in\mathcal{A}(G)$ be such that
@@
As applications of Theorem 1.1, we present two corollaries, which demonstrate that almost all (in a combinatorial sense) length sets with maximal elasticity are intervals. To establish the first result, we require a lemma concerning the growth behavior of the $n$-fold sumset of a finite set. Khvanskii first proved that the $n$-fold sumset of a finite set $A\subset\mathbb{Z}^{d}$ grows like a polynomial. His result was generalized and refined, and the bounds were made effective by many authors ([15, 10, 5]). We cite a version due to Granville, Shakan, and Walker ([10]).
@@
We have $\mathcal{A}(\mathcal{B}_{\rho}(G))\subset\mathcal{B}_{\rho}(G)\subset\mathcal{F}(G)$. Set $\Omega:=\psi(\mathcal{A}(\mathcal{B}_{\rho}(G)))$. In particular, the monoid $\mathcal{B}_{\rho}(G)$ is isomorphic to the monoid $S:=[\Omega]$ generated by $\Omega$.
@@
Let $\alpha_{N}\in\mathbb{N}$ be such that $\alpha_{N}N\geq\alpha$. We claim that
@@
Hence the claim follows. Next, since
$|\{B\in\mathcal{B}_{\rho}(G):|B|\leq n-\alpha\}|\geq|\{B\in\mathcal{B}_{\rho}(G):|B|\leq n-\alpha_{N}N\}|,$
it suffices to show that
@@
Let $G=C_{n_{1}}\oplus\ldots\oplus C_{n_{r}}$ with $r\in\mathbb{N}_{0}$ and $1<n_{1}\,|\,\ldots\,|\,n_{r}$, and set $\mathsf{D}^{*}(G)=1+\sum_{i=1}^{r}(n_{i}-1)$. It only takes a few lines to check that
@@
    $|U_{1}|=|U_{2}|=\mathsf{D}(G),\hskip 2.84544ptg_{1}g_{2}\hskip 2.84544pt|\hskip 2.84544ptU_{1},\hskip 5.69046pt\text{and }\hskip 2.84544pt(g_{1}+g_{2})\hskip 2.84544pt|\hskip 2.84544ptU_{2}\,$
is satisfied in each of the following cases:
@@
    $G$ is a cyclic group of odd order
@@
    $G$ is not cyclic and $\mathsf{D}(G)=\mathsf{D}^{*}(G)$
@@
    $G$ has odd order and there is some $U\in\mathcal{A}(G)$ of length $|U|=\mathsf{D}(G)$ that is not squarefree
@@
(a) Suppose $G$ is cyclic of odd order $n\geq 3$. Let $g\in G$ be any generator. Clearly, $2g\in G$ is also a generator of the group, since the order of $G$ is odd. Now, consider $U_{1}=g^{n}$ and $U_{2}=(2g)^{n}$. Since both $U_{1}$ and $U_{2}$ are atoms with length $\mathsf{D}(G)$, our condition holds.
(b) Suppose that $G\cong C_{n_{1}}\oplus\ldots\oplus C_{n_{r}}$ with $1<n_{1}\mid\ldots\mid n_{r}$ and that $\mathsf{D}^{*}(G)=\mathsf{D}(G)$. Let $(e_{1},\ldots,e_{r})$ be a basis of $G$ with $\operatorname{ord}(e_{i})=n_{i}$ for $i\in[1,r]$ and with $r\geq 2$. Then the sequence $U=(e_{1}+e_{2}+\ldots+e_{r})\prod_{i=1}^{r}e_{i}^{n_{i}-1}$ has the property that $|U|=\mathsf{D}(G)$. Moreover, the map $\varphi:G\longrightarrow G$, given by $e_{i}\mapsto e_{i}$ for all $i\in[1,r-1]$ and $e_{r}\mapsto e_{r}+e_{r-1}$, is an isomorphism. Since $U$ is a minimal zero-sum sequence, $\varphi(U)$ is as well, and therefore, $|\varphi(U)|=|U|=\mathsf{D}(G)$ and $e_{r-1}+e_{r}\in\textup{supp }(\varphi(U))$. If we choose $U_{1}=U$ and $U_{2}=\varphi(U)$, then the condition holds.
(c) Suppose $G$ has odd order and there is some $U\in\mathcal{A}(G)$ of length $|U|=\mathsf{D}(G)$ that is not squarefree. Then the map defined by the multiplication by two, $\varphi:G\longrightarrow G,\hskip 5.69046ptg\mapsto 2g$ for every $g\in G$, is an isomorphism. Since $U$ is not squarefree, there exists $g\in G$ such that $g^{2}\hskip 2.84544pt|\hskip 2.84544ptU$. Then $\varphi(U)$ is a minimal zero-sum sequence of length $|\varphi(U)|=|U|=\mathsf{D}(G)$ and $2g\in\textup{supp }(\varphi(U))$. Thus, the condition holds with $U_{1}=U$ and $U_{2}=\varphi(U)$.
@@
We end this section by considering a Property P of finite abelian groups and its arithmetic consequences. Theorem 4.3 establishes a result on the catenary degree for elements with maximal elasticity. It is analogous to the result given by Proposition 3.1, but its validity is quite restricted. Let $H$ be an atomic monoid and let $a\in H$. Then the basic inequality (2.1) shows that $\mathsf{c}(a)\leq 3$ implies that $\mathsf{L}(a)$ is an interval. But the converse is far from being true. Thus, the conclusion of Theorem 4.3 is stronger than that of Theorem 1.1.
@@
Let $H$ be a transfer Krull monoid over a finite abelian group $G$, and suppose that $G$ satisfies Property P. Then there exists some $a^{*}\in H$ such that all elements $a\in H$ with $\rho(\mathsf{L}(a))=\rho(H)$ have catenary degree $\mathsf{c}(aa^{*})\leq 3$, whence the length sets $\mathsf{L}(a^{*}a)$ are intervals with elasticity $\rho(H)$.
@@
where the sequences $A_{g}$ are as in the Definition of Property P. Lemma 3.3 follows that the length set of $A^{*}$ has maximal elasticity. Let $A\in\mathcal{B}(G)$ be such that $\rho(\mathsf{L}(A))=\mathsf{D}(G)/2$. Then $\mathsf{\rho}(\mathsf{L}(A^{*}A))=\mathsf{D}(G)/2$ by Lemma 3.2. Since $\textup{supp }(A^{*})\cup\{0\}=G$, we get $\textup{supp }(A^{*}A)\cup\{0\}=G$. Thus, [6, Theorem 7.6.8] implies that $\mathsf{c}(A^{*}A)\leq 3$, and consequently, the length set $\mathsf{L}(A^{*}A)$ forms an interval. It has elasticity $\rho(H)$ by Lemma 3.2. ∎